: A firm produces three products in a repetitive process facility. Product A sells for $100; its variable costs are $20. Product B sells for $250; its variable costs are $90. Product C sells for $40; its variable costs are $10. The firm has annual fixed costs of $350,000. Last year, the firm sold 800 units of A, 1800 units of B, and 5,000 units of C. Calculate the break-even point of the firm. The firm has some idle capacity at these volumes, and chooses to cut the selling price of A from $100 to $50, believing that its sales volume will rise from 800 units to 2000 units. What is the revised break-even point?
Answer #1

Break-even point of the firm (in units) = Total Fixed Costs

Contribution per unit

For Product A : Fixed Costs = $ 350,000

Contribution = Selling price per unit - variable cost per unit

= $ 100 - $ 20 = $ 80

For Product B : Fixed Costs = $ 350,000

Contribution = Selling price per unit - variable cost per unit

= $ 250 - $ 90 = $ 160

For Product C : Fixed Costs = $ 350,000

Contribution = Selling price per unit - variable cost per unit

= $ 40 - $ 10 = $ 30

Break Even Point Product A Product B Product C
$350,000/80 = 4375 units $350,000/160 = 2187.5 units $350,000/30 = 11666.67 units

Revised Break even point

For Product A :

Fixed Costs = $ 350,000

Contribution = Selling price per unit - variable cost per unit

= $ 50 - $ 20 = $ 30

Break even point = fixed costs/Contribution per unit

= $ 350,000/30 = $ 11666.67

Here, we can say that selling price for first 800 units will be $ 100 and for next 1200 units selling price will be $ 50.